The amount of workdone in pumping water out of a cubical vessel of ht 1m is

Grade:12th pass

1 Answers

Arun
5 years ago
Work = Change in Potential energy
= Final Potential energy - Initial Potential energy
= mgL - mgL/2
= mgL/2
= (ALd gL)/2
= (1^2 * 1 * 10^3 * 10 * 1)/2
= 5 kJ